• what is the mass in grams of 1.204 x 10²³ molecules of H2O?


• how many formula units are there in 0.175 g of LiF

1
Expert's answer
2022-02-02T04:12:04-0500

Solution:

(1):

One mole of any substance contains 6.022×1023 particles.

Therefore,

Moles of H2O = (1.204×1023 molecules) × (1 mol / 6.022×1023 molecules) = 0.1999 mol = 0.2 mol


The molar mass of H2O is 18.0153 g/mol

Therefore,

Mass of H2O = (0.2 mol H2O) × (18.0153 g H2O / 1 mol H2O) = 3.603 g H2O = 3.6 g H2O


Answer (1): 3.6 grams of H2O


(2):

The molar mass of LiF is 25.94 g/mol

Therefore,

Moles of LiF = (0.175 g LiF) × (1 mol LiF / 25.94 g LiF) = 0.00675 mol LiF


One mole of any substance contains 6.022×1023 particles.

Therefore,

Formula units of LiF = (0.00675 mol) × (6.022×1023 formula units / 1 mol) = 4.065×1021 formula units


Answer (2): 4.065×1021 formula units LiF